An experimenter is convinced that her measuring instrument had a variability measured by variance σ2 = 4. During an experiement she recorded the measurements 4.1, 5.2, and 10.2. What is the value of the test statistic to test the hypotheses H0 : σ2 = 4 vs H1 : σ2 6= 4

A. 3.212

B. 5.285

C. 4.111

D. 8.34
Answer #1

Here given values are

n=3

sigma^2 =4

values=4.1,5.2,10.2

mean =(4.1+5.2+10.2)/3=19.5/3=6.5

sample variance = s^2 =(1/2)*(((4.1-6.5)^2)+((5.2-6.5)^2)+(10.2-6.5)^2) =21.14/2 =10.57

then here we use chisquare test, the test statistic is given by

chi square value =((n-1)*(s^2))/(sigma^2)

= ((3-1)*10.57)/4 = 21.14 / 4 =5.285

so option B

if you are satisfied with my answer then please give me a thumbsup.

please drop a thumbs up

thank you, have a nice day.